OpenStudy (anonymous):

simplify square root of 5 x square root of 6 please. keep getting lost

OpenStudy (anonymous):

\[\sqrt{5}*\sqrt{6}\] \[=\sqrt{30}\]

OpenStudy (anonymous):

ok , \[\sqrt{5}*\sqrt{6} \] is this the question right

OpenStudy (anonymous):

\[\sqrt{30}\]

OpenStudy (anonymous):

we an right this as whole square root of 5*6 that is 30 \[\sqrt{30}\]

OpenStudy (anonymous):

yes that is the right question

OpenStudy (anonymous):

\[\sqrt{5}\sqrt{6}=\sqrt{30}\]

OpenStudy (anonymous):

or u can say we have take square root as common . got it:)

OpenStudy (anonymous):

is this the simplest form

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes : since if we divide 30 into factors , no one can be put out hence it is a simplest form